# Representative Action Selection for Large Action Space Meta-Bandits

### v2

@@ Appendix D Proof of Theorem 8
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 7,
@@
where the first equality uses tower rule. The first inequality uses Equations (12-13), and Lemma 17. The last equality uses tower rule again. The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
Note that the Gaussian process assumption is not used in Equation (14); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. By applying Equation (10) in the proof of Theorem 7, we bound this term and complete the proof. ∎
Appendix E Proof of Theorem 10
@@
Let =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
@@
which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over [0,1)[0,1)[ 0 , 1 ) at
@@
Case II $m\geq K+1$: Consider the relaxed maximization problem of Equation (15):
@@
To find critical points of Equation (16), we solve the system:
@@
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of Equation (16). The corresponding objective value is:
@@
Finally, note that Equation (16) is a relaxation of Equation (15). While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of Equation (15) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
@@
Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$. Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
@@
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 16 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
@@
where the first inequality follows from the algorithm-dependent upper bound in Theorem 8. The second inequality follows from Equations (17) and (18), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 18). The last inequality comes from
@@
where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$. Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $-\tfrac{K}{2K+1}\approx-\frac{1}{2}$. The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of Vershynin [2018]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix F Proof of Theorem 11
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 6. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 7 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
@@ Appendix F Proof of Theorem 11
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 17. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [Dragomir, 2004], corresponding to a constant $c>0$:
@@
Appendix G Proof of Theorem 12
@@
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [Talagrand, 2014] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
@@
Let $\mathcal{A}$ be output of Algorithm 1. Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Then, for the same constant $C>0$,
@@
The first inequality follows from Equation (14) in the proof of Theorem 8. The last inequality uses Lemma 20. ∎
Appendix H The effect of clustering structure on regret
Appendix I Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [Williams and Rasmussen, 2006].
[…]
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (20), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$. In this way, we sample functions from a RKHS function class; See [Kanagawa et al., 2018, Theorem 4.12].
Figure 3: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (20). Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in [0,2][0,2][ 0 , 2 ] (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval [0,2][0,2][ 0 , 2 ], using Gibbs kernel in Equation (20) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. Figure 3 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
Appendix J Independent and identically distributed actions
As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d. random variables. This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of 111 at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
Appendix K Properties of Gaussian width
@@ Appendix I Varying-dependence actions with RBF/Gibbs kernels
where the first equality uses $-\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $-a\in-\mathcal{S}$.
@@
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need this inequality to prove the next claim.
@@
Appendix L Bounds of expected maximum of Gaussian
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [Kamath, 2015]:

### v3

@@ Appendix E Proof of Theorem 7
The key tool for decomposing the regret of Algorithm 1 into that from a reference subset and that from missing clusters is the following Lemma:
@@
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 6,
@@
where the first equality uses tower rule. The first inequality uses Equations (11-12), and Lemma 16. The last equality uses tower rule again. The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
Note that the Gaussian process assumption is not used in equation 13; it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. By applying equation 9 in the proof of Theorem 6, we bound this term and complete the proof. ∎
Appendix F Proof of Theorem 9
To analyze the worst-case behavior of Theorem 7, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
@@ Appendix F Proof of Theorem 9
Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
@@
which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over $[0,1)$ at
@@
Case II $m\geq K+1$: Consider the relaxed maximization problem of equation 14:
@@
To find critical points of equation 15, we solve the system:
@@
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of equation 15. The corresponding objective value is:
@@
Finally, note that equation 15 is a relaxation of equation 14. While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of equation 14 is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
@@
Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$. Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
@@
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 15 to shift; and the final bound uses Claim 3 from Supplementary I. Then,
@@
where the first inequality follows from the algorithm-dependent upper bound in Theorem 7. The second inequality follows from Equations equation 16 and equation 17, and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 17). The last inequality comes from
@@
where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$. Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $\tfrac{K}{2K+1}\approx-\frac{1}{2}$. The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [37]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix G Proof of Theorem 10
To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma 16:
Statement of Theorem 10
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 5. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 6 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
@@ Appendix G Proof of Theorem 10
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 16. The second inequality follows from equation 18. The third inequality uses the Pólya-Szegő inequality [11], corresponding to a constant $c>0$:
@@
Appendix H Proof of Theorem 11
@@
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [34] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
@@
Let $\mathcal{A}$ be output of Algorithm 1. Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Then, for the same constant $C>0$,
@@
The first inequality follows from equation 13 in the proof of Theorem 7. The last inequality uses Lemma 19. ∎
Appendix I Properties of Gaussian width
@@
where the first equality uses $\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $a\in-\mathcal{S}$.
@@
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need inequality to prove the next claim.
@@
Appendix J Bounds of expected maximum of Gaussian
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [17]:
@@
Appendix K The effect of clustering structure on regret
Appendix L Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [39].
$\begin{split}\underset{\text{RBF}}{k}(a,a^{\prime})=&\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right),\\ \underset{\text{Gibbs}}{k}(a,a^{\prime})=&\sqrt{\frac{2\,l(a)l(a^{\prime})}{l(a)^{2}\!+\!l(a^{\prime})^{2}}}\exp\left(\!-\!\frac{\|a-a^{\prime}\|^{2}}{l(a)^{2}\!+\!l(a^{\prime})^{2}}\right),\end{split}$ (19)
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in equation 19, the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$. In this way, we sample functions from a RKHS function class; See [18, Theorem 4.12].
Figure 4: Experiments with outcome functions sampled from RBF/Gibbs kernels in equation 19. Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in equation 19 to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. Figure 4 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
Appendix M Independent and identically distributed actions
As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d. random variables. This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of $1$ at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary J, we have
Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in equation 1,